OpenStudy (anonymous):

Solve the following system of equations. 2x + y = 3 x = 2y - 1

OpenStudy (anonymous):

Plug in the second equation into the first and solve for x then with x you can use the second equation to solve for y.. this is similar to the one i worked out for you once before except the second equation is already solved for x

OpenStudy (anonymous):

@jewjewbri what did you get for your answer? i think the answer for both x and y are the same.

OpenStudy (anonymous):

2x and 2y?

OpenStudy (anonymous):

This is what i did. 2x + y =3 2 ( 2y - 1) + y = 3 4y -2 + y = 3 5y = 5 y = 1

OpenStudy (anonymous):

now you have the value of y which is 1 try solving for x now